# references.bib

### v4

@@ 3.2 Index theorem for H-elliptic operators from $KK$-theoretic viewpoint
Finally, we conclude the index classes and symbol classes in different cases above, and lead to a basic expectation for transversally H-elliptic operators. The construction of symbol and index classes in $KK$-homology, and their correlation to the analytical and topological index of the operator on a compact manifold are summarized as follows:
@@
  • The index class, expressed as the $KK$-pairing $[(L^{2}(E),P)]\in KK_{G}(A,\mathbb{C})$, with different preference of the $C^{\ast}$-algebra $A$ in varying cases. The index class represents the analytical index for operators.
@@
  • The symbol class is represented as the $KK$-pairing $[(C(M;E)\otimes_{C(M)}B,\sigma(P))]\in KK_{G}(\mathbb{C},B)$, where $B$ is selected to be the symbol algebra of “negative order” terms in different cases. It is constructed from the symbol depending on the topological properties of $P$ and $M$, such symbol class can be understood as the topological index. Especially, in the transversally elliptic case, the “negative order” algebra $S_{tr}(M)$ is replaced by the $KK$-equivalent Clifford algebra $Cl_{\Gamma}(TM)=C_{0}(TM)\otimes_{C(M)}Cl_{\Gamma}(M)$.
@@
In this chapter, we explains the definition of transversally H-elliptic operators.
@@ 4.1 Basic notions
Let $M$ be a compact filtered Riemannian manifold equipped with a smooth, isometric action of a Lie group $G$. Recall that for each point $x\in M$, we have the mapping $f_{x}:G\to M$, $g\mapsto g(x)$. The derivative of $f_{x}$ at $e\in G$ is denoted by $f_{x}^{\prime}:\mathfrak{g}\to T_{x}M$.
@@
and we obtain a $G$-invariant map $\tilde{\varphi}_{x}=h_{x}h_{x}^{\ast}:gr(H)^{\ast}_{x}\to gr(H)^{\ast}_{x}$.
@@
To define transversally H-elliptic operators from the perspective of the tangent Clifford symbol class, we first define the space of Clifford sections $Cl_{\tilde{\Gamma}}(M)$. Recall that we have defined $Cl_{V}(M)$ in Definition 2.2.4 for any vector bundle $V$.
Definition 4.1.4.
Define the vector bundle
$\tilde{\Gamma}:=\{h(s)\ |\ s\in C_{0}(M;M\times\mathfrak{g})\}.$
The map $f^{\prime}$ in Definition 2.2.4 is replaced by the $h$ map.
Then $Cl_{\tilde{\Gamma}}(M)$ is the algebra of continuous sections of the sub-algebra of $\text{Cliff}(gr(H))$ generated by the vectors tangent to the orbit direction.
@@ 4.2 Construction of leaf-wise Dirac operator
  • Then we construct the Dirac operator acting on sections of the vector bundle $\Lambda^{\ast}(\mathcal{O}_{x})$ for each orbit $\mathcal{O}_{x}\cong G/M_{x}$. Using the homeomorphisms $f_{x}:G/{M_{x}}\cong\mathcal{O}_{x}$, we can push forward all these operators $\mathcal{D}_{G/M_{x}}$ to the orbits $\mathcal{O}_{x}$ in $M$. The operator $\mathcal{D}$ on the orbit $O$ will be denoted $\mathcal{D}_{O}$. By the construction above, it depends only on the orbit, not on the point $x$ of the orbit.
  • Using the embedding $\mathfrak{m}^{\bot}_{x}\cong T_{x}(O_{x})\subset T_{x}(M)$, one can redefine the operators $\mathcal{D}_{O_{x}}$ as operators acting on the bundles $\Lambda^{\ast}(M)|_{O_{x}}$. This will give the operators $\mathcal{D}_{O}$ on sections of the bundle $\Lambda^{\ast}(M)$ over the orbit $O$.
@@
    The orbital Dirac operator $\mathcal{D}_{G}$ on $\Lambda^{\ast}(M)$ is formed by the family of operators $\{\mathcal{D}_{O_{x}}\}$ on $\Lambda^{\ast}(M)|_{O_{x}}$.
@@
There exist an orbital Dirac operator $\mathcal{D}_{G}$ on $\Lambda^{\ast}(M)$ whose ordinary principle symbol is $\text{ext}(\varphi_{x}(\xi))+\text{int}(\varphi_{x}(\xi))$.
@@
The existence has been verified already. The ordinary principle symbol can be calculated by
$\sigma({\mathcal{D}_{G}})(x,\xi)$ $=\sum_{k}(\xi,f_{x}^{\prime}(v_{k}))(\text{ext}(f_{x}^{\prime}(v_{k}))+\text{int}(f_{x}^{\prime}(v_{k})))$
$=\sum_{k}({f_{x}^{\prime}}^{\ast}(\xi),v_{k})(\text{ext}(f_{x}^{\prime}(v_{k}))+\text{int}(f_{x}^{\prime}(v_{k})))$
$=\text{ext}(\varphi_{x}(\xi))+\text{int}(\varphi_{x}(\xi)).$
@@ 4.3 Adjoint in different viewpoints
The H-principle cosymbol $\sigma_{H}(\mathcal{D}_{G},x)$ at one point $x$ should be an operator on $\mathcal{G}=Gr(H)_{x}$. It defines an operator on $C_{c}^{\infty}(\mathcal{G})$ and extends to $C^{\ast}(\mathcal{G})$. The main purpose of this and the next section is to show the essential self-adjointness of
@@
Note that we can define the adjoint operator of a left-invariant differential operator $P=\sigma_{H}(\mathcal{D}_{G},x)$, for example, in the sense of $C^{\ast}$-module morphism or in the sense of universal algebra. However, it turns out that these adjoint maps $P\to P^{\ast}$ are in fact the same.
@@ 4.4 Essential self-adjointness of $\sigma_{H}(\mathcal{D}_{G})$
Let $\mathcal{D}_{G}$ be the orbital Dirac operator on $\Lambda^{\ast}(M)$. For each $x\in M$, take the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$. Then the H-principle cosymbol of $\mathcal{D}_{G}$ is an unbounded operator on the Hilbert module $E$
@@
In this section, we want to prove the following theorem:
@@
the limit is $w$.
By Mazur’s lemma, there is a sequence $(v_{n})$ in $C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ consisting of convex combinations of the sequence, such that $v_{n}\to v$ and $F_{n}v\to w$ which strongly converges to $w$ as $n\to\infty$. Thus we have proved that when $v$ in the maximal domain is compactly supported, $v$ is in the minimal domain of $D$.
@@ 4.5 Definition for transversally H-elliptic operators
From the basic idea of transversally elliptic theory, we shall define the transversal H-ellipticity by requiring operators to be H-elliptic in the directions transversal to $G$-orbits.
The fundamental definition is motivated by the tangent-Clifford symbol for transversally elliptic operators defined in Definition 2.2.5. Recall that for a transversally elliptic operator $P$, its tangent-Clifford symbol $[\sigma^{tcl}(P)]$ is in the $KK$-group $KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))=KK_{G}(\mathbb{C},Cl_{\Gamma}(M)\otimes_{C(M)}C_{0}(TM))$.
Replacing the $C_{0}(TM)\cong C^{\ast}(TM)$ by $C^{\ast}(Gr(H))$ and $Cl_{{\Gamma}}(M)$ by $Cl_{\tilde{\Gamma}}(M)$ in the group $KK_{G}(\mathbb{C},Cl_{\Gamma}(TM))=KK_{G}(\mathbb{C},Cl_{\Gamma}(M)\otimes_{C(M)}C_{0}(TM))$, we intend to derive the formal formula of the tangent-Clifford symbol for operators that can induce a $KK$-class $[\sigma_{H}^{tcl}(P)]$ in
$KK_{G}(\mathbb{C},Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H))).$
The definition of $KK$-classes requires $1-\sigma^{tcl}_{H}(P)^{2}$ to be a compact multiplier, which motivates our definition.
@@
Let $M$ be a compact filtered Riemannian manifold and $G$ be a Lie group which acts on $M$ isometrically and preserves the filtered structure. We call a pseudo-differential operator of 00-order $P$ is transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in M(E_{x}\otimes_{\mathbb{C}}C^{\ast}(Gr(H)_{x}))$
@@
  • In the multiplier term $(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}$, the first term $(1-\sigma_{H}^{2}(P,x))$ is a multiplier of $E_{x}\otimes C^{\ast}(Gr(H)_{x})$ by definition of H-cosymbol, and the second term $(1+\sigma_{H}(\Delta_{G},x))^{-1}$, by the essentially self-adjointness of $\sigma_{H}(\mathcal{D}_{G},x)$ as a $C^{\ast}(Gr(H)_{x})$ morphism in Theorem 4.4.1, also represent a multiplier of $C^{\ast}(Gr(H)_{x})$. When vector bundle $E$ is trivial (in general cases we take the local trivialization such that $\text{End}(E)$ part becomes a matrix), the condition in Definition 4.5.1 requires that for all $x\in M$, the product multiplier element $(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}$ lies in the $C^{\ast}$-algebra $C^{\ast}(Gr(H)_{x})$.
  • The leaf-wise sub-Laplacian operator $\Delta_{G}$ is employed instead of a leaf-wise H-elliptic operator, because the leaf-wise direction contains several different order terms if an additional estimate is not obtained. Thus, an appropriate operator like $\Delta_{G}$ with H-elliptic property cannot be readily presumed in the leaf-wise direction in general cases.
@@
If the group $G$ acts trivially on $M$, then the condition will reduce to $1-\sigma_{H}^{2}(P,x)$ is compact. In this case, $1-\sigma_{H}^{2}(P,x)$ is homogeneous up to smooth so in the pseudo-differential calculus, the compactness of a multiplier is equivalent to of negative order. The condition is equivalent to the existence of parametrix for $\sigma(P,x)$, which implies the Rockland condition at each point $x\in M$ and leads to the H-ellipticity, Conversely, it can be proved that the cosymbols of H-elliptic operators are unbounded Fredholm multipliers (refer to [joachim2003unbounded] for the definition). Considering the bounded transform expressed in Section 3.2, then $1-\sigma^{2}_{H}(P)$ is a compact multiplier.
@@
Define the operator $\sigma_{H}^{tcl}(P)$ by
$\sigma^{tcl}_{H}(P):=\sigma_{H}(P)\hat{\otimes}1+(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f_{\tilde{\Gamma}}\in\mathcal{L}(C(M;E)\hat{\otimes}_{C(M)}C^{\ast}(Gr(H))\otimes_{C(M)}Cl_{\tilde{\Gamma}}(M)).$
Here $f_{\tilde{\Gamma}}$ is a multiplier of $Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H))$ defined as $\sigma_{H}(\mathcal{D}_{G})\otimes(1+\sigma_{H}(\Delta_{G}))^{-1/2}$. At each point, $\sigma_{H}(\mathcal{D}_{G},x)$ denotes an element in Clifford product space $\tilde{\Gamma}_{x}$ such that it represents a multiplier on $Cl_{\tilde{\Gamma}}(M)$, $(1+\sigma_{H}(\Delta_{G,x}))^{-1/2}$ is a multiplier on $C^{\ast}(Gr(H)_{x})$. Consider the tensor product over $C(M)$ to obtain the desired operator $f_{\tilde{\Gamma}}$.
By the construction of $\sigma_{H}^{tcl}(P)$, we have the following
@@
$1-(\sigma^{tcl}_{H}(P))^{2}=(1-\sigma^{2}_{H}(P))(1-f_{\tilde{\Gamma}}^{2})=(1-\sigma_{H}^{2}(P))(1+\sigma_{H}(\Delta_{G}))^{-1}.$
@@
$1-(\sigma^{tcl}_{H}(P))^{2}=1-\sigma^{2}_{H}(P)\hat{\otimes}1-(1-\sigma_{H}(P)^{2})\hat{\otimes}f^{2}_{\tilde{\Gamma}}-(1-\sigma_{H}(P)^{2})^{1/2}[\sigma_{H}(P)\hat{\otimes}1,1\hat{\otimes}f_{\tilde{\Gamma}}].$
Since $\sigma_{H}(P)\hat{\otimes}1$ represents on $C^{\ast}(Gr(H))$ part and $1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})$ represents on $Cl_{\tilde{\Gamma}}(M)$ part, the commutator $[\sigma_{H}(P)\hat{{\otimes}}1,1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})]=0$. Let $f(a)={a}/{(1+a^{2})}$, then $1\hat{\otimes}f_{\tilde{\Gamma}}$ is the functional calculus by $f$ on $1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})$. As a result, $[\sigma_{H}(P)\hat{\otimes}1,1\hat{\otimes}f_{\tilde{\Gamma}}]=0$ and
$1-(\sigma^{tcl}_{H}(P))^{2}=$ $1-\sigma^{2}_{H}(P)\hat{\otimes}1-(1-\sigma_{H}(P)^{2})\hat{\otimes}f^{2}_{\tilde{\Gamma}}$
$=$ $(1-\sigma^{2}_{H}(P))(1-f_{\tilde{\Gamma}}^{2})$
@@
Thus $1-(\sigma_{H}^{tcl}(P))^{2}$ is compact by definition of transversal H-ellipticity. So we can define that:
@@
Let $P$ be a transversally H-elliptic operator. The symbol class $[\sigma_{H}^{tcl}(P)]\in K^{G}_{0}(Cl_{\tilde{\Gamma}}(M)\otimes_{C(M)}C^{\ast}(Gr(H)))$ is defined by the $KK$-pairing
$[\sigma_{H}^{tcl}(P)]:=(C(M;E)\hat{\otimes}_{C(M)}C^{\ast}(Gr(H)){\otimes}Cl_{\tilde{\Gamma}}(M),\sigma^{tcl}_{H}(P)).$
@@
Let $P$ be a transversally H-elliptic operator. Then $(L^{2}(E),P)$ defines an element in the group $K^{0}(C^{\ast}(G))$. It is denoted $[P]$ and called the index class of $P$.

### v5

@@ 3.3 Index theorem for H-elliptic operators from $KK$-theoretic viewpoint
Finally, we conclude the index classes and symbol classes in different cases, which lead to a basic expectation for transversally H-elliptic operators. The construction of symbol and index classes in $KK$-homology, and their correlation to the analytical and topological index of the operator on a locally compact manifold are summarized as follows:
@@
  • The index class, expressed as the $KK$-pairing $[(L^{2}(E),P)]\in KK^{G}(A,\mathbb{C})$, with different preference of the $C^{\ast}$-algebra $A$ in varying cases. The index class represents the analytical index for operators.
@@
  • The symbol class is represented as the $KK$-pairing $[(C_{0}(M;E)\otimes_{C_{0}(M)}B,\sigma(P))]\in KK^{G}(C_{0}(M),B)$, where $B$ is selected to be the symbol algebra of “negative order” terms in different cases. It is constructed from the symbol depending on the topological properties of $P$ and $M$, such symbol class can be understood as the topological index. Especially, in the transversally elliptic case, the “negative order” algebra $S_{tr}(M)$ is replaced by the $KK$-equivalent Clifford algebra $Cl_{\Gamma}(TM)=C_{0}(TM)\otimes_{C_{0}(M)}Cl_{\Gamma}(M)$.
@@
From now on, we begin to construct the definition of transversally H-elliptic operators from H-elliptic theory.
@@ 4.1 Basic notions
Let $M$ be a complete filtered Riemannian manifold equipped with a smooth, proper, and isometric action of a locally compact Lie group $G$. Recall that for each point $x\in M$, we have the mapping $f_{x}:G\to M$, $g\mapsto g(x)$. The derivative of $f_{x}$ at $e\in G$ is denoted by $f_{x}^{\prime}:\mathfrak{g}\to T_{x}M$.
To transform the concepts in transversal elliptic theory to transversal H-elliptic theory, we need to construct a natural isomorphism between the tangent bundle $TM$ and the bundle of osculating Lie algebras $gr(H)=\bigsqcup_{x}gr(H)_{x}$.
@@
and we obtain a $G$-invariant map $\tilde{\varphi}_{x}=(h_{x}h_{x}^{\ast})^{1/2}:gr(H)^{\ast}_{x}\to gr(H)^{\ast}_{x}$.
@@ 4.2 Construction of leaf-wise Dirac operator
The difference between H-elliptic and transversally H-elliptic operators lies in the leaf-wise part of the $G$-action. So we need a leaf-wise elliptic operator, the orbital Dirac operator, so that it will intertwine with transversally H-elliptic operators in some sense to become H-elliptic operators.
@@
    Proof.
    By direct calculation, the principal symbol of $\mathcal{D}_{G/M_{x}}$ at $\eta$ is
    $-i\sum_{k}i\langle\eta,v_{k}\rangle(\text{ext}(\tilde{v}_{k})+\text{int}(\tilde{v}_{k}))=\text{ext}(\eta)+\text{int}(\eta).$
    It is invertible for aribitrary $\eta\neq 0$, thus $\mathcal{D}_{G/M_{x}}$ is elliptic. ∎
  • Then we construct the Dirac operator acting on sections of the vector bundle $\Lambda^{\ast}(\mathcal{O}_{x})$ for each orbit $\mathcal{O}_{x}\cong G/M_{x}$. Using the homeomorphisms $f_{x}:G/{M_{x}}\cong\mathcal{O}_{x}$, we can push forward all these operators $\mathcal{D}_{G/M_{x}}$ to the orbits $\mathcal{O}_{x}$ in $M$. The operator $\mathcal{D}$ on the orbit $\mathcal{O}$ will be denoted $\mathcal{D}_{\mathcal{O}}$. By the construction above, it depends only on the orbit, not on the point $x$ of the orbit.
  • Using the embedding $\mathfrak{m}^{\bot}_{x}\cong T_{x}({\mathcal{O}}_{x})\subset T_{x}(M)$, one can redefine the operators $\mathcal{D}_{{\mathcal{O}}_{x}}$ as operators acting on the bundles $\Lambda^{\ast}(M)|_{{\mathcal{O}}_{x}}$. This will give the operators $\mathcal{D}_{\mathcal{O}}$ on sections of the bundle $\Lambda^{\ast}(M)$ over the orbit ${\mathcal{O}}$.
@@
    The orbital Dirac operator $\mathcal{D}_{G}$ on $\Lambda^{\ast}(M)$ is formed by the family of operators $\{\mathcal{D}_{{\mathcal{O}}_{x}}\}$ on $\Lambda^{\ast}(M)|_{{\mathcal{O}}_{x}}$.
@@
There exist an orbital Dirac operator $\mathcal{D}_{G}$ on $\Lambda^{\ast}(M)$ whose ordinary principal symbol is ext ($\text{ext }(\varphi_{x}(\xi))+\text{int }(\varphi_{x}(\xi))$.
@@
The existence has been verified already. The ordinary principal symbol can be calculated as follows. By Lemma 4.2.4, the principal symbol of $D_{\mathcal{O}}$ on $\Lambda^{\ast}(M)|_{{\mathcal{O}}}$ can be written as
$\sigma({D_{\mathcal{O}}})(x,\xi)=\text{ext}({f^{\prime}_{x}}^{\ast}(\xi))+\text{int}({f^{\prime}_{x}}^{\ast}(\xi)).$
Note that ${f^{\prime}_{x}}^{\ast}(\xi)$ lies in $\mathfrak{g}_{x}^{\ast}$, and we want to transform it back to $T_{x}M$. Since $f^{\prime}_{x}$ is a linear map from $\mathfrak{g}$ to $T_{x}M$, the image of the linear map $f^{\prime}_{x}$ and the image of its adjoint ${f^{\prime}_{x}}^{\ast}$ are isometrically isomorphic by
$(f^{\prime}_{x}{f^{\prime}_{x}}^{\ast})^{-1/2}f^{\prime}_{x}:\text{Im}(f^{\prime}_{x})\to\text{Im}({f^{\prime}_{x}}^{\ast}).$
Applying this isometry to ${f^{\prime}_{x}}^{\ast}$, we can replace it by
$(f^{\prime}_{x}{f^{\prime}_{x}}^{\ast})^{-1/2}f^{\prime}_{x}=(f^{\prime}_{x}{f^{\prime}_{x}}^{\ast})^{1/2}(\xi)=\varphi_{x}(\xi)$
which is defined in Section 2.2. Therefore, we have
$\sigma(\mathcal{D}_{G})(x,\xi)=\text{ext}(\varphi_{x}(\xi))+\text{int}(\varphi_{x}(\xi)).$
@@ 4.3 Adjoint in different viewpoints
The H-principal cosymbol $\sigma_{H}(\mathcal{D}_{G},x)$ at one point $x$ should be an operator on $\mathcal{G}=Gr(H)_{x}$. It defines an operator on $C_{c}^{\infty}(\mathcal{G})$ and extends to $C^{\ast}(\mathcal{G})$. In order to give a well-defined definition for transversally H-elliptic operators (Definition 4.5.1), it is reasonable to show the essential self-adjointness of
@@
We can define the adjoint operator of a left-invariant differential operator $P=\sigma_{H}(\mathcal{D}_{G},x)$, for example, in the sense of $C^{\ast}$-module morphism or in the sense of universal algebra. Before we verify the essential self-adjointness of $\sigma_{H}(\mathcal{D}_{G},x)$, it suffices to show that these adjoint maps $P\to P^{\ast}$ are in fact the same.
@@ 4.4 Essential self-adjointness of $\sigma_{H}(\mathcal{D}_{G})$
In this section, we show the essential self-adjointness of $\sigma_{H}(\mathcal{D}_{G})$, where $\mathcal{D}_{G}$ is the orbital Dirac operator on $\Lambda^{\ast}(M)$ constructed in Section 4.2.
For each $x\in M$, take the $C^{\ast}(Gr(H)_{x})$-module $E=C^{\ast}(Gr(H)_{x})\otimes\Lambda^{\ast}_{x}(M)$. Then the H-principal cosymbol of $\mathcal{D}_{G}$ is an unbounded operator on the Hilbert module $E$
@@
In this section, we will prove the following theorem:
@@
the limit is $w$. Here we need the following lemma:
Lemma 4.4.3 (Mazur’s lemma).
Let $(X,\lVert\cdot\rVert)$ be a normed vector space and let $\{x_{j}\}_{j\in\mathbb{N}}\subset X$ be a sequence converges weakly to some $x\in X$. Then there exists a sequence $\{y_{k}\}_{k\in\mathbb{N}}\subset X$ made up of finite convex combination of the $x_{j}$’s of the form
$y_{k}=\sum_{j\geq k}\lambda_{j}^{(k)}x_{j}$
such that $y_{k}$ is strongly convergent to $x$, i.e., $\lVert y_{k}-x\rVert\to 0$ as $k\to\infty$.
By Lemma 4.4.3, there is a sequence $(v_{n})$ in $C_{c}^{\infty}({Gr(H)_{x}})\otimes{\Lambda^{\ast}_{x}(M)}$ consisting of convex combinations of the sequence, such that $v_{n}\to v$ and $F_{n}v\to w$ which strongly converges to $w$ as $n\to\infty$. Thus we have proved that when $v$ in the maximal domain is compactly supported, $v$ is in the minimal domain of $D$.
@@ 4.5 Definition for transversally H-elliptic operators
Now we can give the definition for transversally H-elliptic operators. From the basic idea of transversally elliptic theory, we shall define the transversal H-ellipticity by requiring operators to be H-elliptic after intertwining with the leafwise elliptic operator $\Delta_{G}$.
The fundamental definition is motivated by the tangent-Clifford symbol for transversally elliptic operators defined in Definition 2.2.5. Recall that for a transversally elliptic operator $P$, its tangent-Clifford symbol $[\sigma^{tcl}(P)]$ is in the $KK$-group $KK^{G}(C_{0}(M),Cl_{\Gamma}(TM))=KK^{G}(C_{0}(M),Cl_{\Gamma}(M)\otimes_{C_{0}(M)}C_{0}(TM))$.
Replacing $C_{0}(TM)\cong C^{\ast}(TM)$ by $C^{\ast}(Gr(H))$ in the group $KK^{G}(C_{0}(M),Cl_{\Gamma}(TM))=KK^{G}(C_{0}(M),Cl_{\Gamma}(M)\otimes_{C_{0}(M)}C_{0}(TM))$, we intend to derive the formal formula of the tangent-Clifford symbol for operators that can induce a $KK$-class $[\sigma_{H}^{tcl}(P)]$ in
$KK^{G}(C_{0}(M),Cl_{{\Gamma}}(M)\hat{\otimes}_{C_{0}(M)}C^{\ast}(Gr(H))).$
The definition of $KK$-classes requires $1-\sigma^{tcl}_{H}(P)^{2}$ to be a compact multiplier, which motivates the definition for transversally H-elliptic operators:
@@
Let $M$ be a complete filtered Riemannian manifold and $G$ be a locally compact Lie group which acts on $M$ properly, isometrically, and preserves the filtered structure. We call a pseudo-differential operator of 00-order $P$ is transversally H-elliptic if at each point $x\in M$, the multiplier term
$(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}\in\mathcal{M}(E_{x}\otimes C^{\ast}(Gr(H)_{x}))$
@@
  • In the multiplier term $(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}$, the first term $(1-\sigma_{H}^{2}(P,x))$ is a multiplier of $E_{x}\otimes C^{\ast}(Gr(H)_{x})$ by definition of H-cosymbol, and the second term $(1+\sigma_{H}(\Delta_{G},x))^{-1}$, by the essentially self-adjointness of $\sigma_{H}(\mathcal{D}_{G},x)$ as a $C^{\ast}(Gr(H)_{x})$ morphism in Theorem 4.4.1, also represent a multiplier of $C^{\ast}(Gr(H)_{x})$. When the vector bundle $E$ is trivial (in general cases we take the local trivialization such that $\text{End}(E)$ part becomes a matrix), the condition in Definition 4.5.1 requires that for all $x\in M$, the product multiplier element $(1-\sigma_{H}^{2}(P,x))(1+\sigma_{H}(\Delta_{G},x))^{-1}$ lies in the $C^{\ast}$-algebra $C^{\ast}(Gr(H)_{x})$.
  • The leaf-wise sub-Laplacian operator $\Delta_{G}$ is employed instead of a leaf-wise H-elliptic operator, because the leaf-wise direction contains several different order terms if no additional conditions are taken. Thus, an appropriate operator like $\Delta_{G}$ with H-elliptic property in the orbit direction cannot be readily presumed in the leaf-wise direction in general cases.
@@
If the group $G$ acts trivially on $M$, then the condition will be reduced to $1-\sigma_{H}^{2}(P,x)$ is compact. In this case, $1-\sigma_{H}^{2}(P,x)$ is homogeneous up to smooth so in the pseudo-differential calculus, the compactness of a multiplier is equivalent to of negative order. The condition is equivalent to the existence of parametrix for $\sigma_{H}(P,x)$, which implies the Rockland condition at each point $x\in M$ and leads to the H-ellipticity, Conversely, it can be proved that the cosymbols of H-elliptic operators are unbounded Fredholm multipliers (refer to [joachim2003unbounded] for the definition). Considering the bounded transform expressed in Section 3.2, then $1-\sigma^{2}_{H}(P)$ is a compact multiplier.
@@
We define the tangent-Clifford symbol $\sigma_{H}^{tcl}(P)$ by
$\sigma^{tcl}_{H}(P):=\sigma_{H}(P)\hat{\otimes}1+(1-\sigma_{H}(P)^{2})^{1/2}\hat{\otimes}f^{\prime}_{{\Gamma}}\in\mathcal{L}(C_{0}(M;E)\hat{\otimes}_{C_{0}(M)}C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M)).$
Here $f^{\prime}_{{\Gamma}}$ is a multiplier of $C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M)$ defined as $\sigma_{H}(\mathcal{D}_{G})\hat{\otimes}(1+\sigma_{H}(\Delta_{G}))^{-1/2}$. By the discussion of $\sigma_{H}(\mathcal{D}_{G})$ in Section 4.4, at each point $x$, $\sigma_{H}(\mathcal{D}_{G},x)$ consists of the sum of the tensor product of an element in $gr(H)_{x}$, which represents a multiplier on $C^{\ast}(Gr(H)_{x})$, and a Clifford-product element on $\Lambda^{\ast}(T_{x}M)$. The Clifford-product part is leaf-wise, so it represents a multiplier on $Cl_{{\Gamma_{x}}}(M)$. The remaining part $(1+\sigma_{H}(\Delta_{G},x))^{-1/2}$ is a multiplier on $C^{\ast}(Gr(H)_{x})$. As a result, we can take $f^{\prime}_{\Gamma}=\sigma_{H}(\mathcal{D}_{G})\hat{\otimes}(1+\sigma_{H}(\Delta_{G}))^{-1/2}$ as a multiplier on $C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M)$.
By the construction of $\sigma_{H}^{tcl}(P)$, we see that
@@
Modulo compact operators on $Gr(H)_{x}$ for each $x$, we have
$1-(\sigma^{tcl}_{H}(P))^{2}=(1-\sigma^{2}_{H}(P))(1-{f^{\prime}_{\Gamma}}^{2})=(1-\sigma_{H}^{2}(P))(1+\sigma_{H}(\Delta_{G}))^{-1}$
@@
$1-(\sigma^{tcl}_{H}(P))^{2}=1-\sigma^{2}_{H}(P)\hat{\otimes}1-(1-\sigma_{H}(P)^{2})\hat{\otimes}{f^{\prime}_{\Gamma}}^{2}-(1-\sigma_{H}(P)^{2})^{1/2}[\sigma_{H}(P)\hat{\otimes}1,1\hat{\otimes}f^{\prime}_{\Gamma}].$
Since the cosymbol $\sigma_{H}(P)\hat{\otimes}1$ acts on bundle $E$ and $1\hat{\otimes}\sigma_{H}(\mathcal{D}_{G})$ represents on $Cl_{{\Gamma}}(M)$, the commutator $[\sigma_{H}(P)\hat{\otimes}1,1\hat{\otimes}f^{\prime}_{\Gamma}]=0$ modulo negative order terms, i.e. modulo compact operators on $Gr(H)_{x}$ for each $x\in M$ because all these operators are left-invariant on $Gr(H)_{x}$. So, modulo compact operators, we have
$1-(\sigma^{tcl}_{H}(P))^{2}=$ $1-\sigma^{2}_{H}(P)\hat{\otimes}1-(1-\sigma_{H}(P)^{2})\hat{\otimes}{f^{\prime}_{\Gamma}}^{2}$
$=$ $(1-\sigma^{2}_{H}(P))(1-{f^{\prime}_{\Gamma}}^{2})$
@@
Thus, by Definition 4.5.1, $1-(\sigma_{H}^{tcl}(P))^{2}$ is compact for transversally H-elliptic operators. So we can define that:
@@
Let $P$ be a transversally H-elliptic operator. The tangent-Clifford cosymbol class $[\sigma_{H}^{tcl}(P)]\in KK^{G}(C_{0}(M),C^{\ast}(Gr(H))\hat{\otimes}_{C_{0}(M)}Cl_{{\Gamma}}(M))$ is defined by the $KK$-pairing
$[\sigma_{H}^{tcl}(P)]:=(C_{0}(M;E)\hat{\otimes}_{C_{0}(M)}C^{\ast}(Gr(H)){\hat{\otimes}_{C_{0}(M)}}Cl_{{\Gamma}}(M),\sigma^{tcl}_{H}(P)).$
@@
Let $P$ be a transversally H-elliptic operator. Then $(L^{2}(E),P)$ defines an element in the group $K^{0}(C^{\ast}(G,C_{0}(M))$. It is denoted $[P]$ and called the index class of $P$.